# Thermodynamics of Schwarzschild-AdS black hole in non-commutative geometry

### v5

@@ III Thermodynamics
$G=M-TS.$ (51)
@@
$\mathrm{d}G=\left(W^{-1}-1\right)T\mathrm{d}S+W^{-1}V\mathrm{d}P-S\mathrm{d}T.$ (56)
@@
Figure 10: At the critical points, the first-order phase transitions occur. The slope of the Gibbs free energy with respect to temperature diverges to infinity. The red points are the critical points.
@@ IV Joule-Thomson process
Figure 12: The constant mass curves (blue lines) and the inversion curves (red lines) for different parameters $a$. The constant mass curves expand outward as the mass increases. We set (a) $a=0.5,M=0.51,M_{i}^{min}\left(\simeq 0.52\right),0.53,0.54,0.55$; (b) $a=1,M=1.02,M_{i}^{min}\left(\simeq 1.04\right),1.06,1.08,1.1$; (c) $a=1.5,M=1.52,M_{i}^{min}\left(\simeq 1.56\right),1.6,1.64,1.68$; (d) $a=2,M=2.02,M_{i}^{min}\left(\simeq 2.08\right),2.14,2.2,2.26$.

### v6

@@ III Thermodynamics
Now, we investigate the energy of gravitational field using the Komar integral [42, 43]. A timelike Killing vector field in spacetime is $\xi=\partial_{t}$, $\xi^{\mu}=\left(1,0,0,0\right)$. The conserved current is constructed as
$J^{\mu}=R^{\mu\nu}\xi_{\nu}.$ (26)
The corresponding conserved charge is the Komar energy
$\frac{1}{4\pi}\int_{\Sigma}\mathrm{d}^{3}x\sqrt{h}n_{\mu}J^{\nu},$ (27)
where $\Sigma$ is three-dimensional hypersurface with the unit normal vector $n_{\mu}=\left(-\sqrt{f\left(r\right)},0,0,0\right)$, and $h$ is the determinant of the induced metric on the surface $\Sigma$. Additionally, it should be noted that $\nabla_{\mu}\nabla_{\sigma}\xi^{\mu}=R_{\sigma\nu}\xi^{\nu}$, so Eq. 27 can be rewritten as
$\frac{1}{4\pi}\int_{\Sigma}\mathrm{d}^{3}x\sqrt{h}n_{\mu}\nabla_{\nu}\left(\nabla^{\mu}\xi^{\nu}\right).$ (28)
Using Stokes’ theorem, the above equation can be written as
$\frac{1}{4\pi}\int_{\partial\Sigma}\mathrm{d}^{2}x\sqrt{\sigma}n_{\mu}r_{\nu}\nabla^{\mu}\xi^{\nu},$ (29)
where $\partial\Sigma$ denotes the boundary of $\Sigma$ and $\sigma$ is the determinant of the induced metric on the two-dimensional hypersurface $\partial\Sigma$. $r_{\mu}=\left(0,1/\sqrt{f\left(r\right)},0,0\right)$ is the unit normal vector of $\partial\Sigma$. We consider the condition $\Lambda=0$. In this case, one could choose $\partial\Sigma$ to be the two-dimensional sphere at infinity $S^{2}_{\infty}$, then one have a reduced metric $\sigma_{ij}=r^{2}\left(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\right)$. One could obtain
$\frac{1}{4\pi}\int_{S^{2}_{\infty}}\mathrm{d}^{2}x\sqrt{\sigma}n_{\mu}r_{\nu}\nabla^{\mu}\xi^{\nu}$ (30)
$=$ $-\frac{1}{4\pi}\int_{S^{2}_{\infty}}\mathrm{d}^{2}x\sqrt{\sigma}\nabla^{t}\xi^{r}$
$=$ $\lim_{r\to\infty}\left(M-\frac{aM}{r}\right)$
$=$ $M.$
Thus, even due to the Smarr relation of the black hole, as shown in Eq. 23, the total energy of the gravitational field can still be understood as $M$. However, since there is no universally accepted method for determining the energy of the gravitational field, this issue holds further research value. The Komar integral can also be used to examine the Smarr relation. To achieve this task, one can consider the quasi-local energy of the gravitational field when $\Lambda$ doesn’t vanish. Let us consider the following integral
$\frac{1}{4\pi}\int_{S^{2}_{r_{h}}}\mathrm{d}^{2}x\sqrt{\sigma}n_{\mu}r_{\nu}\nabla^{\mu}\xi^{\nu}$ (31)
$=$ $W^{\prime}M-\frac{\Lambda r_{h}^{3}}{3}$
$=$ $W^{\prime}M+2PV,$
where $W^{\prime}=1-\frac{a}{r_{h}}$ is another correction function. And please notice that $W^{\prime}\neq W$. Considering that in static sphericallyy spacetime, one could prove
$\nabla^{t}\xi^{r}\big{|}_{\mathcal{N}}=-\kappa,$ (32)
where $\mathcal{N}$ is Killing horizon and $\kappa$ is black hole’s surface gravity. Additionally, due to $\kappa$ is a constant, one could derive
$\frac{1}{4\pi}\int\mathrm{d}^{2}x\sqrt{\sigma}n_{\mu}r_{\nu}\nabla^{\mu}\xi^{\nu}$ (33)
$=$ $\frac{1}{4\pi}\int\mathrm{d}^{2}x\sqrt{\sigma}\kappa$
$=$ $\frac{\kappa A}{4\pi}$
$=$ $2TS.$
Eq. 31 and Eq. 33 gives
$W^{\prime}M=2TS-2PV.$ (34)
This formula also could be understood as the Smarr relation. This relation is also correct. However, compared to the Smarr relation obtained through a scale transformation, it lacks some physical information because it does not consider the non-commutative parameter as a thermodynamic variable.
@@
$G=M-TS=\frac{r_{h}\left(3r_{h}-8\pi Pr_{h}^{3}+a\left(3+16\pi Pr_{h}^{2}\right)\right)}{6\left(2r_{h}-a\right)}.$ (60)
@@
$\mathrm{d}G=\left(W^{-1}-1\right)T\mathrm{d}S+W^{-1}V\mathrm{d}P-S\mathrm{d}T+W^{-1}\Phi_{a}\mathrm{d}a.$ (65)
@@
Figure 10: At the critical points, the first-order phase transitions occur. The slope of the Gibbs free energy with respect to temperature diverges to infinity. The black points are the critical points.
@@ IV Joule-Thomson process
Figure 12: The constant mass curves (blue solid lines) and the inversion curves (red dashed lines) for different parameters $a$. The constant mass curves expand outward as the mass increases. We set (a) $a=0.5,M=0.51,M_{i}^{min}\left(\simeq 0.52\right),0.53,0.54,0.55$; (b) $a=1,M=1.02,M_{i}^{min}\left(\simeq 1.04\right),1.06,1.08,1.1$; (c) $a=1.5,M=1.52,M_{i}^{min}\left(\simeq 1.56\right),1.6,1.64,1.68$; (d) $a=2,M=2.02,M_{i}^{min}\left(\simeq 2.08\right),2.14,2.2,2.26$.